An equalateral triangle and a circle have the same center. The area of triangle not in the circle = the area of the circle not in the triangle. If the radius of the circle is 1 find the length of the side of the triangle.
Answer by venugopalramana(3286)

An equalateral triangle and a circle have the same center. The area of triangle not in the circle = the area of the circle not in the triangle. If the radius of the circle is 1 find the length of the side of the triangle.
WE HAVE FROM THE GIVEN DIPOITION OF THE FIGURE
AREA OF EQ.TRIANGLE-AREA OF CIRCLE+AREA OF CIRCLE OUTSIDE TRIANGLE=AREA OF EQ.TRIANGLE OUTSIDE CIRCLE
BUT AREA OF EQ.TRIANGLE OUTSIDE CIRCLE=AREA OF CIRCLE OUTSIDE EQ.TRIANGLE
AREA OF EQ.TRIANGLE-AREA OF CIRCLE =0
HENCE AREA OF EQ.TRIANGLE=AREA OF CIRCLE
IF SIDE OF EQ.TRIANGLE =B ...THEN
B^2*SQRT(3)/4=PI*1*1=3.14
B^2=4*3.14/SQRT(3)=4*3.14/1.732
B=2.69